OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. Please show all of your work. 5/6-{-7/12}

OpenStudy (anonymous):

= 5/6 + 7/12 = 10/12 + 7/12 = 17/12 or 1 5/12

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

tw